# ES-GNN: Generalizing Graph Neural Networks Beyond Homophily with Edge Splitting

### v3

@@ 3.1.2 Connection to GNNs Considering Task-Relevance
The concept of prioritizing task-relevance in GNNs has been extensively explored across various domains, such as topological denoising [48, 49, 50], graph pooling [51], augmentations [52, 53], contrastive learning [54], and structure learning [55, 56, 57]. Given our focus on supervised classification tasks, the forthcoming discussion will be expressly centered around this area. This focus ensures a contextual analysis of ES-GNN within the established research landscape, highlighting its unique contributions to task-oriented GNN development.
@@ 3.2 Disentangled Representation Learning
Disentangled representation learning, aimed at disentangling the explanatory latent variables within observed data into distinct dimensions [58, 59], has garnered considerable attention, particularly in the field of computer vision [60, 61, 62, 63, 64]. In recent years, there has been a progressive expansion of disentangled representation learning into the graph domain, addressing a wide spectrum of tasks. These range from foundational classification [65, 66, 33, 67, 68, 69, 70, 71, 72, 73, 74, 75] and generation challenges [76, 77, 78, 79, 80], in both supervised and unsupervised settings, to downstream applications like trajectory prediction [81], recommendation systems [82, 83, 84, 85, 86], and graph neural architecture search [87]. Given our focus on foundational Graph Neural Network (GNN) models, the following discussion will concentrate on the subset of research that employs disentangled representation learning to enhance the capabilities of GNNs within the realm of fundamental tasks.
@@ 4.2 Aggregation Layer
4.3 Irrelevant Consistency Regularization
Stacking ES-layer and aggregation layer iteratively lends itself to disentangling different features of nodes into two distinct representations, denoted by $\mathbf{Z}_{\text{\text{R}}}$ and $\mathbf{Z}_{\text{\text{IR}}}$. First, $\mathbf{Z}_{\text{R}}$, informed and shaped by $\mathbf{A}_{\text{R}}$, is tuned for prediction, with its development guided by the minimization of the classification loss $\mathcal{L}_{\text{pred}}$. This process not only makes $\mathbf{Z}_{\text{R}}$ predictive of node labels but also implicitly reinforces the task-relevant nature of $\mathbf{A}_{\text{R}}$ via message passing. However, only supervising one channel (R) risks neglecting the meaningfulness of the other (IR), potentially leading to the preservation of erroneous information in predictions. To this end, we introduce a Irrelevant Consistency Regularization (ICR) loss $\mathcal{L}_{\text{ICR}}$, designed to regulate $\mathbf{Z}_{\text{IR}}$ as the opposite of $\mathbf{Z}_{\text{R}}$, i.e., identifying the classification-harmful information within the observed graph.
The key rationale is to explore the similarities among nodes that are detrimental to classification tasks within $\mathbf{Z}_{\text{\text{IR}}}$. Given any node pairs $(v_{i},v_{j})\in\mathcal{E}$, we would expect $\mathbf{Z}_{{\text{\text{IR}}}[i,:]}$ and $\mathbf{Z}_{{\text{\text{IR}}}[j,:]}$ to be close in the latent space if they possess different labels. Specifically, our ICR can be formulated as:
where $\delta$ is a Kronecker function returning 1 if $\mathbf{y}_{i}=\mathbf{y}_{j}$ and 0 otherwise, and $\|\cdot\|_{2}$ denotes $L_{2}$ norm. As such, $\mathbf{Z}_{{\text{\text{IR}}}}$ is constrained with a local consistency between adjacent nodes from different classes. As a benefit, the classification-harmful information between nodes can be further excluded from task-relevant features $\mathbf{Z}_{\text{\text{R}}}$, and disentangled in the task-irrelevant ones $\mathbf{Z}_{\text{\text{IR}}}$.
Several powerful techniques [90, 27] have been developed to measure the label-agreement between pairwise nodes. In this work, however, we find that using directly the joint probability from model prediction works well, which also offers advantages in low computational complexity as no additional trainable parameters are required.
@@ 5.2 Disentangled Smoothness Assumption in ES-GNN
$\mathop{\arg\min}_{\mathbf{Z}_{\text{R}},\mathbf{Z}_{\text{IR}}}$ $\|\mathbf{Z}_{\text{R}}-{\color[rgb]{0,0,0}{\mathbf{X}_{\text{R}}}}\|^{2}_{2}+\|\mathbf{Z}_{\text{\text{IR}}}-\mathbf{X}_{\text{\text{IR}}}\|^{2}_{2}$ (12)
$+\xi\cdot tr(\mathbf{Z}^{T}_{\text{R}}\mathbf{L}_{\text{R}}\mathbf{Z}_{\text{R}})+\xi\cdot tr(\mathbf{Z}^{T}_{\text{IR}}\mathbf{L}_{\text{IR}}\mathbf{Z}_{\text{IR}})$
@@
$\mathop{\arg\min}_{\mathbf{Z}_{\text{\text{R}}}^{*}}\|\mathbf{Z}_{\text{\text{R}}}^{*}-\mathbf{Z}_{\text{\text{R}}}^{(0)}\|^{2}_{2}+\xi\cdot tr({\mathbf{Z}_{\text{\text{R}}}^{*}}^{T}\mathbf{L}_{\text{\text{R}}}^{(k)}\mathbf{Z}_{\text{\text{R}}}^{*})$ (15)
$\mathop{\arg\min}_{\mathbf{Z}_{\text{\text{IR}}}^{*}}\|\mathbf{Z}_{\text{\text{IR}}}^{*}-\mathbf{Z}_{\text{\text{IR}}}^{(0)}\|^{2}_{2}+\xi\cdot tr({\mathbf{Z}_{\text{\text{IR}}}^{*}}^{T}\mathbf{L}_{\text{\text{\text{IR}}}}^{(k)}\mathbf{Z}_{\text{\text{IR}}}^{*})$ (16)
@@
When adopting the normalized Laplacian matrix $\mathbf{L}_{\text{\text{R}}}=\mathbf{I}-\mathbf{D}_{\text{\text{R}}}^{-\frac{1}{2}}\mathbf{A}_{\text{\text{R}}}\mathbf{D}_{\text{\text{R}}}^{-\frac{1}{2}}$, the feature aggregation operator in Eq. (7) with channel $s=\text{\text{R}}$ can be regarded as solving Eq. (15) using iterative gradient descent with stepsize $\beta=\frac{1}{2+2\xi}$ and $\xi=\frac{1}{\epsilon_{\text{\text{R}}}}-1$.
@@
$=2\beta\mathbf{Z}_{\text{R}}^{(0)}+2\beta\xi(\mathbf{D}_{\text{R}}^{-\frac{1}{2}}\mathbf{A}_{\text{R}}\mathbf{D}_{\text{R}}^{-\frac{1}{2}})\mathbf{Z}_{\text{R}}^{(k)}+(1-2\beta-2\beta\xi)\mathbf{Z}_{\text{R}}^{(k)}.$ (18)
Setting $\beta$ as $\frac{1}{2+2\xi}$ gives us:
$\mathbf{Z}_{\text{R}}^{(k+1)}=\frac{1}{1+\xi}\mathbf{Z}_{\text{R}}^{(0)}+\frac{\xi}{1+\xi}(\mathbf{D}_{\text{R}}^{-\frac{1}{2}}\mathbf{A}_{\text{R}}\mathbf{D}_{\text{R}}^{-\frac{1}{2}})\mathbf{Z}_{\text{R}}^{(k)},$ (19)
which is equivalent to Eq. (7) while choosing $\xi=\frac{1}{\epsilon_{\text{R}}}-1$, i.e.,
@@ 6.1.5 Implementation Details
For all the baselines and our model, we set $d=64$ as the number of hidden states for fair comparison, and tune the hyper-parameters on the validation split of each dataset using Optuna [99] for 200 trials. With the best hyper-parameters, we train models in 1,000 epochs using the early-stopping strategy with a patience of 100 epochs. We then report the models’ average performance across 10 runs on the test set for each of the 10 random splits, leading to a total of 100 runs. For reproducibility, we provide the searching space of our hyper-parameters: learning rate $\sim\left[1\mathrm{e}{-2},1\mathrm{e}{-1}\right]$, weight decay $\sim\left[1\mathrm{e}{-6},1\mathrm{e}{-3}\right]$, dropout $\sim\{0,0.1,...,0.8\}$ with step $0.1$, the number of layers $K\sim\{1,2,...,8\}$ with step 1, scaling parameter $\epsilon_{\text{R}},\epsilon_{\text{IR}}\sim\{0.1,0.2,...,1\}$ with step $0.1$, and irrelevant consistency coefficient $\lambda_{\text{ICR}}\sim\left[0,1\right]$ for Cora, Citeseer, Pubmed, and Twitch-DE, $\left[5\mathrm{e}{-8},5\mathrm{e}{-6}\right]$ for Chameleon, Wisconsin, Cornell, and Texas, $\left[5\mathrm{e}{-5},5\mathrm{e}{-3}\right]$ for Squirrel, and $\left[5\mathrm{e}{-3},5\mathrm{e}{-2}\right]$ for Actor. Our implementation can be found at https://github.com/jingweio/ES-GNN.
@@ 6.6 Robustness Analysis
(a) Squirrel
(b) Chameleon
(c) Twitch-DE
(d) Actor
(e) Cora
(f) Citeseer
(g) Pubmed
(h) Polblogs
Figure 9: Sensitivity analysis of coefficient $\lambda_{\text{ICR}}$.
@@ 6.8 Channel Analysis and Ablation Study
6.9 Sensitivity Analysis of Coefficient $\lambda_{\text{ICR}}$
We test the effect of the irrelevant consistency coefficient $\lambda_{\text{ICR}}$, and plot the learning performance of our model on eight real-world datasets as examples in Fig. 9 by varying $\lambda_{\text{ICR}}$ with different values. For example, the classification accuracy on Squirrel in Fig. 9a goes up first and then gradually drops. Promising results can be attained by choosing $\lambda_{\text{ICR}}$ from $\left[5\mathrm{e}{-5},5\mathrm{e}{-3}\right]$. Similar trends can be also observed on the other datasets, where $\lambda_{\text{ICR}}$ is relatively robust within a wide albeit distinct interval.
@@ 7 Conclusion
Acknowledgments
The work was partially supported by the following: National Natural Science Foundation of China under no.61876155; Jiangsu Science and Technology Programme (Natural Science Foundation of Jiangsu Province) under no. BK20181189, BE2020006-4; Key Program Special Fund in XJTLU under no. KSF-A-10, KSF-T-06, KSF-E-26, KSF-P-02, and KSF-A-01.

### v5

@@ 3.1.2 Connection to GNNs Considering Task-Relevance
The concept of prioritizing task-relevance in GNNs has been extensively explored across various domains, such as topological denoising [48, 49, 50], graph pooling [51], augmentations [52, 53], contrastive learning [54], and structure learning [55, 56, 57]. Although [58] does not explicitly define task-relevance or irrelevance, it subtly explains the reasons for node connections by employing relational topic modeling with hierarchical graph information. Given our focus on supervised classification tasks, the forthcoming discussion will be expressly centered around this area. This focus ensures a contextual analysis of ES-GNN within the established research landscape, highlighting its unique contributions to task-oriented GNN development.
@@ 3.2 Disentangled Representation Learning
Disentangled representation learning, aimed at disentangling the explanatory latent variables within observed data into distinct dimensions [59, 60], has garnered considerable attention, particularly in the field of computer vision [61, 62, 63, 64, 65]. In recent years, there has been a progressive expansion of disentangled representation learning into the graph domain, addressing a wide spectrum of tasks. These range from foundational classification [66, 67, 33, 68, 69, 70, 71, 72, 73, 74, 75, 76] and generation challenges [77, 78, 79, 80, 81], in both supervised and unsupervised settings, to downstream applications like trajectory prediction [82], overlapping community detection [83], recommendation systems [84, 85, 86, 87, 88], and graph neural architecture search [89]. Given our focus on foundational Graph Neural Network (GNN) models, the following discussion will concentrate on the subset of research that employs disentangled representation learning to enhance the capabilities of GNNs within the realm of fundamental tasks.
@@ 4.2 Aggregation Layer
4.3 Irrelevant Consistency Regularization
Stacking ES-layer and aggregation layer iteratively lends itself to disentangling different features of nodes into two distinct representations, denoted by $\mathbf{Z}_{\text{\text{R}}}$ and $\mathbf{Z}_{\text{\text{IR}}}$. First, $\mathbf{Z}_{\text{R}}$, informed and shaped by $\mathbf{A}_{\text{R}}$, is tuned for prediction, with its development guided by the minimization of the classification loss $\mathcal{L}_{\text{pred}}$. This process not only makes $\mathbf{Z}_{\text{R}}$ predictive of node labels but also implicitly reinforces the task-relevant nature of $\mathbf{A}_{\text{R}}$ via message passing. However, only supervising one channel (R) risks neglecting the meaningfulness of the other (IR), potentially leading to the preservation of erroneous information in predictions. To this end, we introduce a Irrelevant Consistency Regularization (ICR) loss $\mathcal{L}_{\text{ICR}}$, designed to regulate $\mathbf{Z}_{\text{IR}}$ as the opposite of $\mathbf{Z}_{\text{R}}$, i.e., identifying the classification-harmful information within the observed graph. The key rationale is to explore the similarities among nodes that are detrimental to classification tasks within $\mathbf{Z}_{\text{\text{IR}}}$. Given any node pairs $(v_{i},v_{j})\in\mathcal{E}$, we would expect $\mathbf{Z}_{{\text{\text{IR}}}[i,:]}$ and $\mathbf{Z}_{{\text{\text{IR}}}[j,:]}$ to be close in the latent space if they possess different labels, which can be formulated as:
where $\delta$ is a Kronecker function (1 if $\mathbf{y}_{i}=\mathbf{y}_{j}$, 0 otherwise) and $\|\cdot\|_{2}$ denotes $L_{2}$ norm. As such, $\mathbf{Z}_{{\text{\text{IR}}}}$ is constrained with a local consistency between adjacent nodes from different classes, aiding in the exclusion and disentanglement of classification-harmful information from $\mathbf{Z}_{\text{\text{R}}}$ and to $\mathbf{Z}_{\text{\text{IR}}}$.
Several powerful techniques have been developed to assess label agreement between nodes [92, 27]. In this work, however, we find that using the joint probability from model predictions is effective and eliminates the need for additional trainable parameters. Besides, while the idea in ICR could be adapted to task-relevant representations by making closer nodes with the same predicted label, we avoid this due to the risk of inaccurate prediction during training. Such inaccuracy could irreversibly distort the classification metric space. Instead, we supervise task-irrelevant representations – unused for direct prediction – with noisy labels, offering a margin of error that our proposed layers can correct before reaching the final prediction stage.
@@ 5.2 Disentangled Smoothness Assumption in ES-GNN
$\mathop{\arg\min}_{\mathbf{Z}_{\text{R}},\mathbf{Z}_{\text{IR}}}$ $\|\mathbf{Z}_{\text{R}}-\mathbf{X}_{\text{R}}\|^{2}_{2}+\|\mathbf{Z}_{\text{\text{IR}}}-\mathbf{X}_{\text{\text{IR}}}\|^{2}_{2}$ (12)
$+\xi_{\text{R}}\cdot tr(\mathbf{Z}^{T}_{\text{R}}\mathbf{L}_{\text{R}}\mathbf{Z}_{\text{R}})+\xi_{\text{IR}}\cdot tr(\mathbf{Z}^{T}_{\text{IR}}\mathbf{L}_{\text{IR}}\mathbf{Z}_{\text{IR}})$
@@
$\mathop{\arg\min}_{\mathbf{Z}_{\text{\text{R}}}^{*}}\|\mathbf{Z}_{\text{\text{R}}}^{*}-\mathbf{Z}_{\text{\text{R}}}^{(0)}\|^{2}_{2}+\xi_{\text{R}}\cdot tr({\mathbf{Z}_{\text{\text{R}}}^{*}}^{T}\mathbf{L}_{\text{\text{R}}}^{(k)}\mathbf{Z}_{\text{\text{R}}}^{*})$ (15)
$\mathop{\arg\min}_{\mathbf{Z}_{\text{\text{IR}}}^{*}}\|\mathbf{Z}_{\text{\text{IR}}}^{*}-\mathbf{Z}_{\text{\text{IR}}}^{(0)}\|^{2}_{2}+\xi_{\text{IR}}\cdot tr({\mathbf{Z}_{\text{\text{IR}}}^{*}}^{T}\mathbf{L}_{\text{\text{\text{IR}}}}^{(k)}\mathbf{Z}_{\text{\text{IR}}}^{*})$ (16)
@@
When adopting the normalized Laplacian matrix $\mathbf{L}_{\text{\text{R}}}=\mathbf{I}-\mathbf{D}_{\text{\text{R}}}^{-\frac{1}{2}}\mathbf{A}_{\text{\text{R}}}\mathbf{D}_{\text{\text{R}}}^{-\frac{1}{2}}$, the feature aggregation operator in Eq. (7) with channel $s=\text{\text{R}}$ can be regarded as solving Eq. (15) using iterative gradient descent with stepsize $\beta=\frac{1}{2+2\xi_{\text{R}}}$ and $\xi_{\text{R}}=\frac{1}{\epsilon_{\text{\text{R}}}}-1$.
@@
$=2\beta\mathbf{Z}_{\text{R}}^{(0)}+2\beta\xi_{\text{R}}(\mathbf{D}_{\text{R}}^{-\frac{1}{2}}\mathbf{A}_{\text{R}}\mathbf{D}_{\text{R}}^{-\frac{1}{2}})\mathbf{Z}_{\text{R}}^{(k)}+(1-2\beta-2\beta\xi_{\text{R}})\mathbf{Z}_{\text{R}}^{(k)}.$ (18)
Setting $\beta$ as $\frac{1}{2+2\xi_{\text{R}}}$ gives us:
$\mathbf{Z}_{\text{R}}^{(k+1)}=\frac{1}{1+\xi_{\text{R}}}\mathbf{Z}_{\text{R}}^{(0)}+\frac{\xi_{\text{R}}}{1+\xi_{\text{R}}}(\mathbf{D}_{\text{R}}^{-\frac{1}{2}}\mathbf{A}_{\text{R}}\mathbf{D}_{\text{R}}^{-\frac{1}{2}})\mathbf{Z}_{\text{R}}^{(k)},$ (19)
which is equivalent to Eq. (7) while choosing $\xi_{\text{R}}=\frac{1}{\epsilon_{\text{R}}}-1$, i.e.,
@@ 6.1.5 Implementation Details
For all the baselines and our model, we set $d=64$ as the number of hidden states for fair comparison, and tune the hyper-parameters on the validation split of each dataset using Optuna [101] for 200 trials. With the best hyper-parameters, we train models in 1,000 epochs using the early-stopping strategy with a patience of 100 epochs. We then report the models’ average performance across 10 runs on the test set for each of the 10 random splits, leading to a total of 100 runs. For reproducibility, we provide the searching space of our hyper-parameters: learning rate $\sim\left[1\mathrm{e}{-2},1\mathrm{e}{-1}\right]$, weight decay $\sim\left[1\mathrm{e}{-6},1\mathrm{e}{-3}\right]$, dropout $\sim\{0,0.1,...,0.8\}$ with step $0.1$, the number of layers $K\sim\{1,2,...,8\}$ with step 1, scaling parameters $\epsilon_{\text{R}},\epsilon_{\text{IR}}\sim\left[5\mathrm{e}{-2},0.5\right]$, and irrelevant consistency coefficient $\lambda_{\text{ICR}}\sim\left[0,1\right]$ for Cora, Citeseer, Pubmed, and Twitch-DE, $\left[5\mathrm{e}{-8},5\mathrm{e}{-6}\right]$ for Chameleon, Wisconsin, Cornell, and Texas, $\left[5\mathrm{e}{-5},5\mathrm{e}{-3}\right]$ for Squirrel, and $\left[5\mathrm{e}{-3},5\mathrm{e}{-2}\right]$ for Actor. Our implementation can be found at https://github.com/jingweio/ES-GNN.
@@ 6.8 Channel Analysis and Ablation Study
6.9 Parameter Study
This section presents the sensitivity analysis of hyper-parameters, specifically $\lambda_{\text{ICR}}$, $\epsilon_{\text{R}}$ and $\epsilon_{\text{IR}}$, using Chameleon and Cora datasets as typical examples. Fig. 9 illustrates how varying these parameters affect our model’s learning performance. Overall, we have the following observations: 1) The effect of the regularization coefficient $\lambda_{\text{ICR}}$ is depicted in Fig. 9(a)-(b). For instance, the classification accuracy on the Chameleon dataset increases first and then gradually decreases. Favorable results can be attained by choosing $\lambda_{\text{ICR}}$ from $\left[1\mathrm{e}{-7},1\mathrm{e}{-5}\right]$. A similar trend can be also observed on Cora dataset where $\lambda_{\text{ICR}}$ is relatively robust within a wide albeit distinct interval; 2) The influence of the scaling coefficients $\epsilon_{\text{R}}$ and $\epsilon_{\text{IR}}$ is evaluated by adjusting their values from $1\mathrm{e}{-3}$ to $1.0$ on Fig. 9(c)-(d) and Fig. 9(e)-(f), respectively. It can be observed that, despite variations in optimal settings, selecting values between $5\mathrm{e}{-2}$ and $0.5$ consistently yields promising performance.
(a) Chameleon
(b) Cora
(c) Chameleon
(d) Cora
(e) Chameleon
(f) Cora
Figure 9: Sensitivity analysis of hyper-parameters: $\lambda_{\text{ICR}}$, $\epsilon_{\text{R}}$ and $\epsilon_{\text{IR}}$ from top to bottom rows.